Not logged in
Tue Nov 2 23:20:07 2010
List of things to do
Remeasure the cavity response (V/Hz). I thought I had this from when I measured the finesse, but I realized that I did this with the transmitted power and not the error signal, so I don't have the information I want. Once I have this, I can do several things, including:
Convert all the spectra Koji and I measured last week into gyro signal units (rad/s/rHz)
Divide the primary error spectrum by the OLTF of the loop to determine the unsuppressed frequency noise level and compare it with the expected free-running laser noise
Remeasure the PDH box TF using the SR785, then
Put the new plot on the wiki
Refer the measured input-shorted PDH box output noise to the input to determine its contribution to frequency noise, compare this with LISO
Re-confirm the validity of the Tektronix VCO's purported external FM deviation setting. I did this for DC frequency shifts using a calibrator in
, and yesterday I double-checked it by putting the frequency-modulated signal into the RF analyzer to see that the sidebands were at the correct frequencies, but I realized that I was looking around 95 MHz instead of 47.5 MHz, so I was really looking at sidebands on the carrier harmonic. Either way, the frequency spacing was exactly the 1kHz it was supposed to be, but I will make the correct plot and put it here.
Connect the decided-upon signals to CDS so that we can systematically monitor them at any point in the future. Alastair has put the finishing touches on our new-and-improved front end code, so this should be ready to go.
Do a thorough realignment of the cavity in both directions and record the improvement in transmission, then make new measurements of the error signals and OLTF. I was originally going to do this first, but I thought it might be wise to do all the other junk I have done first while Alastair finished up with the front end.